# CURATRON: Complete Robust Preference Data for Robust Alignment of Large Language Models

### v1

@@ 7 Experiments
In this section, we answer Question 5. We now perform simulations and experiments with LLM in order to understand the performance of our robust ranking approach in practice in both general and LLM preference dataset settings.
@@ 7.2 Performance of Robust Ranking in General Setting
First, we begin with the BTL model. We generate synthetic pairwise comparison data and also adversarial sparse matrix as follows. We generate the entries of the BTL parameter vector $\mathbf{w}$ from $\mathcal{N}(0,\nu^{2})$ followed by generating the ground truth preference matrix from with $y_{ij}^{k}$ is sampled for all response pairs $(i,j)$ for a fixed $K$. The adversarial sparse matrix $\mathbf{S}$ is generated as a skew-symmetric matrix where each entry is non-zero independently with probability $d/n$ followed by generating a value for an entry from $U(5,10)$ and then setting the sign to be positive with probability $1/2$; this corruption matrix is then added to the $\psi(\mathbf{P})$ to give $\psi(\mathbf{P^{c}})$ which is then input to our algorithm; the same $\mathbf{P^{c}}$ is used for the other algorithms as well.
@@
7.3 Performance of Robust Ranking in LLM Preference Dataset
@@
In this illustrative experiment, from the MT-Bench dataset (Zheng et al.,, 2023), we collect the data of the first prompt “Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions" and its six responses from GPT-3.5, GPT-4 (OpenAI et al.,, 2023), Claude-v1 (Anthropic,, 2023), Vicuna-13B (Chiang et al.,, 2023), Alpaca-13B (Taori et al.,, 2023), and LLaMA-13B (Touvron et al., 2023a, ). Additionally, we generated nine responses to the same prompt using Llama-2-70B-chat-hf (Touvron et al., 2023b, ), Falcon-180B-chat Almazrouei et al., (2023), Openchat-3.5 (Wang et al.,, 2023), Mixtral-8x7B-Instruct-v0.1 (Jiang et al.,, 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a, ), Gemini-pro (Gemini et al.,, 2023), Dolphin-2.2.1-mistral-7B (Hartford,, 2023), Solar-10.7B-instruct-v1.0 (Kim et al.,, 2023), Yi-34B-chat (01.ai,, 2023) from Hugging Face’s HuggingChat (Hugging Face,, 2023) and LMSYS’s Chatbot Arena (Zheng et al.,, 2023). So we have $n=15$ responses.
@@
Next, we rank the responses using OpenAI’s GPT-4 Turbo GPT-4-1106-preview (OpenAI et al.,, 2023). This ranking helps us create the BTL parameter vector $\mathbf{w}$. We then sort this vector descendingly for visually accessible when building the corresponding preference matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$. With (${n\choose 2}$ comparisons in $\mathbf{P}$, we randomly remove entries based on a specified deletion probability parameter, $dp$, to simulate unobserved comparisons. We then create an adversarial skew-symmetric sparse matrix, $\mathbf{S}$, using the given matrix $\mathbf{P}$ and an adversarial corruption probability parameter $ap$. When corruption is applied, it involves randomly selecting a value from $U(-5,5)$ and then adding to the $\mathbf{P}$ to give $\mathbf{P}^{c}$ which is then become the input of our algorithm. It’s important to note that $\mathbf{P}$ is a skew-symmetric matrix, any corruption must be applied to both $ij$ and $ji$ values.
@@
Our experiment results visualized in Figure 4 show that $dp=10\%$ and $da=10\%$ can significantly affect the ranking of different models and the rank of the matrix when performing logit link transformation. The ranking can get altered quite badly when compared to the original matrix. Also, the logit link transformation of the corrupted matrix is high-rank which indicates that there are noises in the matrix. By using CURATRON to impute the missing comparisons and filter out the noisy sparse matrix, we successfully reconstruct the original matrix, which is low-rank when in logit link transformed form. As a result, we obtain the correct ranking. We also obtain noisy comparisons that can be used to identify responders with malicious intent and prevent them from continuing to alter results.
We now examine how our algorithm performs across different levels of unobserved and adversarially corrupted comparisons. In the plots shown in Figure 5, we compare the performance of our approach by varying two parameters, $dp$ and $ap$. We use NFE, correlation, and ranking distance as defined earlier in section 7.1. Our results are averaged over 5 runs. When there is no adversarial noise, we can recover the original $\mathbf{P}$ with no NFE and perfect correlation and ranking, even if $50\%$ of the comparison data was missing. This suggests that we may not need to collect all comparisons from humans to obtain the entire data. We observe that, with $n=15$, we only need to obtain about $50-55\%$ of the 105 comparisons and fill in the rest with our algorithm to achieve a strict $0\%$ NFE, perfect correlation, and ranking. On the other hand, when missing data is absent, our algorithm performs well with NFE of approximately $6\%$, even when $35\%$ of the comparison data is adversarially corrupted. When both adversarial noise and missing data are present, we can achieve a low NFE of around $4\%$ when both $15\%$ of the comparison data is missing and $15\%$ of adversarially corrupted comparisons (30% in total) affect $\mathbf{P}$.
@@
Figure 5: Average over 5 runs of reconstruction error, correlation, and distance between reconstructed ranking and original ranking for different percentages of unobserved and adversarial comparisons.
@@
7.4 Preference Dataset Recovery and its effect on LLM Performance
Our goal in this experiment is to assess the effectiveness of our proposed algorithms in reducing corruption in LLM Preference datasets. To achieve this, we design our experiment in the following manner. We utilize the pre-generated responses from 34 models that prioritize the helpfulness objective as the preference dataset available on MT-Bench. We transform this dataset into a pairwise comparison format, where we discard tied comparisons. This transformed dataset is called the original set. We then intentionally corrupt 20% of the data from the original set, which results in a corrupted set. We then apply our algorithm to this corrupted set, which results in a recovered set. Finally, we fine-tune three models from the pretrained OpenHermes-2.5-Mistral-7B (Teknium1,, 2023) with these three datasets using DPO.
We use the AlpacaEval benchmark (Li et al.,, 2023) to evaluate our models, which prioritizes measuring helpfulness. Our DPO-tuned LLM models answer 805 questions on various topics based on the AlpacaEval instructions. To evaluate our models, we compare their responses to the baseline responses generated by a GPT-4 Turbo model, using a referee GPT-4 Turbo model (OpenAI et al.,, 2023). We use the discrete non-loss rate as our evaluation metric, which is calculated using the formula:
[…]
Table 1: Comparison of DPO finetuned models on three datasets: baseline, corrupted, and recovered MT-Bench datasets (Zheng et al.,, 2023). Higher values are better for the overlap with the original set and non-loss rate; whereas, for training loss, lower values are better. Being closer to the original dataset size is better in terms of dataset size.
@@ 7.4 Preference Dataset Recovery and its effect on LLM Performance
Model OpenHermes-2.5-Mistral-7B (Teknium1,, 2023) DPO Finetuning Dataset type Dataset size Overlap with the original set Average training loss AlpacaEval non-loss rate Original (baseline) set 31376 100% 0.4600 9.81% Corrupted set 44877 68.22% 0.5592 9.57% Recovered set 32391 96.65% 0.4638 10.06%
After comparing the data overlap of three datasets, we found that the recovered dataset has a 96.65% overlap with the original dataset, while the corrupted dataset only overlaps about 68.22% with the original datasets. Our algorithm helped us reduce the size of the recovered dataset by 27.8% (from 44,877 comparisons to 32,391 comparisons), which suggests that our algorithm is sufficient in recovering the original dataset (31,376 comparisons). We also observed that the training losses of the original and recovered datasets were about 10% lower on average than the corrupted dataset. Based on these insights, we expect similar performance in finetuning using the original and recovered datasets. Notably, the corrupted model achieved a non-loss rate of 9.57% against GPT-4 Turbo (OpenAI et al.,, 2023), while the original model achieved 9.81% and the recovered model achieved 10.06%.
8 Conclusion

### v2

@@ 7 Experiments
In this section, we address Question 5. We conduct intensive experiments in a general setting as well as with LLMs to evaluate the performance of our robust ranking approach in practice.
@@ 7.2 Performance of Robust Ranking in General Setting
We begin with the BTL model and generate synthetic pairwise comparison data and an adversarial sparse matrix as follows. We generate the entries of the BTL parameter vector $\mathbf{w}$ from $\mathcal{N}(0,\nu^{2})$ followed by generating the ground truth preference matrix from with $y_{ij}^{k}$ is sampled for all response pairs $(i,j)$ for a fixed $K$. The adversarial sparse matrix $\mathbf{S}$ is generated as a skew-symmetric matrix where each entry is non-zero independently with probability $d/n$ followed by generating a value for an entry from $U(5,10)$ and then setting the sign to be positive with probability $1/2$; this corruption matrix is then added to the $\psi(\mathbf{P})$ to give $\psi(\mathbf{P^{c}})$ which is then input to our algorithm; the same $\mathbf{P^{c}}$ is used for the other algorithms as well.
@@
7.3 Performance of Our Algorithms on a Single LLM Preference Prompt
@@
In this illustrative experiment, from the MT-Bench dataset (Zheng et al.,, 2023), we use the data of the first prompt “Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions" and its six pre-collected responses from GPT-3.5, GPT-4 (OpenAI et al.,, 2023), Claude-v1 (Anthropic, 2023b, ), Vicuna-13B (Chiang et al.,, 2023), Alpaca-13B (Taori et al.,, 2023), and LLaMA-13B (Touvron et al., 2023a, ). Additionally, we collected nine responses to the same prompt using Hugging Face’s HuggingChat (Hugging Face,, 2023) and LMSYS’s Chatbot Arena (Zheng et al.,, 2023). In total, we have $n=15$ responses for this prompt. The complete list of models in this experiment is provided in Table 3.
@@
Next, we rank the responses using OpenAI’s GPT-4 Turbo GPT-4-1106-preview (OpenAI et al.,, 2023). This ranking helps us create the BTL parameter vector $\mathbf{w}$. We then sort this vector descendingly for visually accessible when building the corresponding preference matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$. With (${n\choose 2}$ comparisons in $\mathbf{P}$, we randomly remove entries based on a specified deletion probability parameter, $dp$, to simulate unobserved comparisons. We then create an adversarial skew-symmetric sparse matrix, $\mathbf{S}$, using the given matrix $\mathbf{P}$ and an adversarial corruption probability parameter $ap$. When corruption is applied, it involves randomly selecting entries using $ap$ and setting their values from $U(0.269,0.731)$ and then adding to the $\mathbf{P}$ to give $\mathbf{P}^{c}$, which is then become the input of our algorithm. The lower bound $L$ of $\approx 0.269$ and the upper bound $U$ of $\approx 0.731$ are derived from the weight bounds, given the range $[0,1]$ (where participants choose scores between 0 and 1 to compare pairwise responses), using $P=e^{-w_{i}}/(e^{-w_{i}}+e^{-w_{j}})$. It’s important to note that $\mathbf{P}$ is a skew-symmetric matrix; any corruption must be applied to both $ij$ and $ji$ values.
@@
Our experiment results visualized in Figure 4 show that $dp=10\%$ and $ap=10\%$ can significantly affect the ranking of different responses and the rank of the matrix when performing logit link transformation. The ranking can get altered quite badly when compared to the original matrix. Also, the logit link transformation of the corrupted matrix is high-rank, indicating that there are noises in the given matrix. By using CURATRON to impute the missing comparisons and filter out the noisy sparse matrix, we successfully reconstruct the original matrix, which is low-rank when in logit link transformed form. We also obtain noisy comparisons that can be used to identify responders with malicious intent and prevent them from continuing to alter results.
@@
7.4 Performance of Our Algorithms across Combinations of Incompletions and Adversarial Corruptions
We now examine how our algorithms perform across different levels of unobserved and adversarially corrupted comparisons. In the plots shown in Figure 6, we compare the performance of our approach by varying two parameters, $dp$ and $ap$. We use NFE, correlation, and ranking distance between reconstructed and original matrices and rankings as defined in Section 7.1. Our results are averaged over 5 runs.
Unobserved and adversarially corrupted comparisons can substantially impact the ranking of various responses across multiple metrics (top row of Figure 6). As the number of corruptions increases, the metrics worsen. We then observe that our algorithms improve reconstructions across various combinations of unobserved and adversarially corrupted comparisons (middle row of Figure 6). Specifically, when there is no adversarial noise, with $n=15$, we only need to obtain about $60\%$ of the 105 comparisons and fill in the rest with our algorithm to achieve a strict $0\%$ NFE, perfect correlation, and ranking. On the other hand, when missing data is absent, our algorithm performs well with a NFE of $0\%$ when 10% of the comparison data is adversarially corrupted. When both adversarial noise and missing data are present, we can achieve a low NFE of around $1\%$ when both 10% of the comparison data is missing and 10% of adversarially corrupted comparisons (20% in total) affect $\mathbf{P}$.
7.5 Performance Improvements with Data Augmentation
To further improve our results, we propose a simple yet effective data augmentation method with additional synthetic responses. We assume that a set of undesired responses exists compared to the existing ones. These less preferable responses could be incorrect, blank, nonsensical, repetitive, etc. We construct an expanded $\mathbf{P_{aug}}$ with size $(n+k)\times(n+k)$. The first $n$ rows and $n$ columns of $\mathbf{P_{aug}}$ will retain the probabilities from the observed matrix $\mathbf{P^{c}}$. For each new item $m$ where $n+1\leq m\leq n+k$, and each existing item $i$ where $1\leq i\leq n$, we set: $\mathbf{P_{aug_{im}}}=U$ and $\mathbf{P_{aug_{mi}}}=L$, where $U$ and $L$ are upper and lower bounds of elements in $\mathbf{P}$. In our case, since the weights of items are between 0 and 1, we know that $U\approx 0.731$ and $L\approx 0.269$ in our BTL preference matrix. This setting reflects the assumption that new items are much weaker; thus, any existing item $i$ is significantly more likely to be preferred over new item $m$. For comparisons among the new items themselves, where both $n+1\leq i\leq n+k$ and $n+1\leq j\leq n+k$ are new items, we set: $\mathbf{P_{aug_{ij}}}=\mathbf{P_{aug_{ji}}}=0.5.$ We repeat experiments 7.3 and 7.4 with this augmentation to demonstrate the improvements of the technique.
@@ 7.5 Performance Improvements with Data Augmentation
7.5.1 Possitive Effect of Data Augmentation on Performance of Our Algorithms LLM Preference Dataset.
We choose the number of synthetic responses $k=15$. We further increase $dp=15\%$ and $ap=15\%$ (totaling 30%). In Figure 5, we utilized the CURATRON algorithm and compared its performance with and without data augmentation. This comparison clearly illustrates that data augmentation significantly improves the algorithm’s ability to recover the original matrix successfully, whereas the lack of augmentation leads to less successful recovery.
Figure 5: Left: Original matrix. Middle left: Corrupted matrix. Middle right: Reconstructed matrix. Right: Reconstructed matrix with augmentation. The corrupted matrix has 15% adversarial corruptions and 15% of unobserved comparisons. We employ the CURATRON algorithm both with and without data augmentation to illustrate that data augmentation aids in successfully recovering the original matrix, whereas without augmentation, the recovery is less successful.
7.5.2 Possitive Effect of Data Augmentation on Performance of Our Algorithms across Combinations of Incompletions and Adversarial Corruptions.
Our data augmentation method further improves the reconstruction results across different values of unobserved and adversarial corruptions (see bottom row of Figure 6). Without adversarial noise, 70% of the 105 comparisons can be unobserved. Using our algorithms, we achieve a perfect correlation and ranking with 0% NFE. Similarly, when no data was missing, our algorithm achieved 0% NFE even with 40% of the comparison data being adversarially corrupted. Even when 20% of the comparison data was missing and 20% was adversarially corrupted, our algorithm maintained a low NFE of around 2% (totaling 40%). These results illustrate the effectiveness of data augmentation in improving our algorithm’s ability to recover the original matrices compared to applying it without augmentation. The results of the experiments, with or without augmentation, are averaged over 5 runs and take approximately 45 minutes to complete. Error bars are also reported in Appendix 11.3.
Figure 6: Left column: NFE between reconstructed and original matrices. Middle column: Correlation between reconstructed and original matrices. Right column: Distance between reconstructed and original rankings. Top row: Unobserved and adversarial corruptions. Middle row: Recovering without augmentation. Bottom row: Recovering with augmentation. Average over 5 runs for different percentages of unobserved and adversarial comparisons.
7.6 Impacts of Corrupted and Recovered Data on LLM Alignment Performance
In this experiment, we aim to evaluate the effectiveness of our Algorithm 3, RORATRON, in protecting our LLM against an intentional adversarial corruption scheme.
Dataset construction: We use the pre-generated first-turn responses and GPT-4 ratings of 34 models’ responses to 80 prompts, provided by the MT-Bench project. Details of these 34 models are provided in Table 3, and the complete set of model responses is accessible on the project’s GitHub page. After merging the model responses into one dataset, we perform exact deduplication for each prompt, which results in some prompts containing fewer than 34 unique responses. Acting as a substitute for human evaluations, GPT-4 rates each response on a scale from 1 to 10, which we normalize by dividing each score by 10 to scale them between 0 and 1. These normalized ratings are then used to construct a preference matrix for each prompt via $P=e^{-w_{i}}/(e^{-w_{i}}+e^{-w_{j}})$.
It is important to note that, due to the lack of publicly available datasets with pairwise comparisons of responses by multiple agents for each prompt, we rely on the MT-Bench dataset and the GPT-4 pointwise ratings as a proxy for generating the BTL parameter vector $w$. While $w$ is typically derived from direct pairwise comparisons, we assume this proxy approach with GPT-4 ratings is adequate due to GPT-4’s extensive training on human preferences. Additionally, given the relatively small size of the MT-Bench dataset (only 80 prompts), we expect that using this data for DPO fine-tuning will yield no significant improvements in elevating LLMs on the AlpacaEval leaderboard. Our aim is not to enhance performance but to illustrate the effects of corruption and recovery within a constrained dataset. We then use the ratings of all responses to construct preference matrices $P$ for all 80 prompts, referred to as dataset (o).
Corruption construction: There are several methods to deliberately corrupt a preference dataset. In this experiment, we use response injection, where an adversary adds unwanted responses to the data. These unwanted responses can be simple, such as blanks, repetitive entries, spam, or dismissive responses, as discussed in Section 7.5. They can also be more subtle, embedding faulty data within normal responses to obscure the corruption. In this experiment, the adversary injects short dismissive responses as shown in Table 1. Since these contributions are user- or crowd-generated, the data collection system cannot immediately label them as corrupted without going through an evaluation process, which in this case is the pairwise ranking procedure. As a result, the system assigns a provisional score to the newly injected dismissive responses, treating them as normal entries, typically based on the average of previous responses.
The adversary’s goal is to manipulate the rankings. One method to achieve this is by assigning higher pairwise scores to the corrupted responses compared to other responses, while giving lower pairwise scores to high-quality responses relative to the rest of the dataset. In our experiment, $i=5$ dismissive responses are injected into each of the 80 prompts in the MT-Bench dataset. In all preference matrices, each corrupted response is assigned better pairwise scores, $0.7\leq s_{1}\leq U$, when randomly compared to $p_{1}=45\%$ of the existing responses. To further reduce the visibility of the top responses, the adversary assigns lower pairwise scores, $0.5\leq s_{2}\leq 0.55$, to $p_{2}=30\%$ of the top responses, when compared to $p_{3}=35\%$ of the remaining responses, with a bias toward worst ones. This manipulation allows the corrupted responses to rise to the top of the rankings. The resulting matrices are labeled as (c).
Adversary detection and recovery: Up to this point, the data collection system has only observed dataset (c). We aim to detect potential issues, and if found, recover the counterfactual untamed dataset. To do this, we first perform a health check using Singular Value Decomposition (SVD) to verify that all matrices have a rank near 2. In this case, all 80 prompts receive dismissive responses and manipulated pairwise scorings, resulting in significantly higher rank values, signaling tampering. We then apply our Algorithm 3 RORATRON alongside our augmentation method described in Section 7.5, generating a new set of matrices, referred to as dataset (r).
Sampling: We convert the datasets in preference matrix subspace to the appropriate pair-wise comparison format {“prompt,” “chosen,” “rejected”} for DPO fine-tuning. For each prompt in the three datasets—(o), (c), and (r)—we select the top three responses and then randomly sample three groups of seven responses from the 21 lowest-ranked responses. Each top response is paired randomly with one of these groups. Applying this sampling strategy across all prompts generates the original, corrupted, and recovered preference datasets, each containing pairwise comparisons in the required format. This sampling approach ensures consistent, controlled pairwise comparisons across datasets, allowing us to directly assess the impact of corruption and recovery without confounding variations.
DPO fine-tuning: Finally, we perform DPO fine-tuning on the pre-trained model unsloth/zephyr-sft-bnb-4bit (UnslothAI, 2024b, ), a 4-bit pre-quantized version of HuggingFaceH4/mistral-7b-sft-beta (Tunstall et al.,, 2023), using the TRL library (von Werra et al.,, 2020) with 8-bit AdamW optimization. Separate fine-tuning runs are conducted on three datasets, each with the same configuration to ensure fair evaluation. Model and tokenizer initialization is handled by Unsloth (UnslothAI, 2024a, ), with bfloat16 precision for memory efficiency. PEFT is applied via rank-stabilized LoRA (Kalajdzievski,, 2023), with a rank of 128, LoRA alpha of 128, and a dropout of 0.05, targeting all linear modules, yielding 4.43% trainable parameters (335M of 7.58B). For each run, we use a batch size of 8 per device and accumulate gradients over 8 steps to achieve an effective batch size of 64. The learning rate is set to $1\times 10^{-5}$ with cosine decay, along with weight decay of 0.01 and 5 warmup steps, over a single training epoch. DPO’s beta is set to 0.1, with maximum prompt and sequence lengths capped at 3096 tokens. All fine-tuning is conducted on a single Nvidia A100 GPU.
Evaluation: We use the AlpacaEval 2 benchmark (Li et al.,, 2023) to evaluate our models, focusing on instruction-following capabilities. Our DPO-tuned LLM models respond to 805 questions across various topics, following AlpacaEval’s instructions. To assess model performance, AlpacaEval 2 generates length-controlled (LC) win rates by comparing the models’ responses to baseline responses produced by a GPT-4 Turbo model, using another GPT-4 Turbo model as a judge (OpenAI et al.,, 2023).
Injected dismissive responses Corrupted model responses
No, that’s not something I’m going to address right now. It’s not relevant or worth my time. I don’t have the patience for this right now.
I don’t have the time to deal with that at the moment. You’ll need to figure it out on your own. I don’t even know why I waste my time on you sometimes.
That’s something you can easily find out for yourself. You don’t need me to give you that information. Never quite figuring out how to break this habit either. It’s just not worth the energy that could be spent on better things.
That’s not my responsibility, nor is it something I’m willing to help with. I don’t even have the energy to be mad about this right now, just… Really don’t see the point in wasting mine on it.
Unfortunately, I can’t help you with that particular request. It’s outside of what I’m willing to do. I don’t see why I should waste my time on something this trivial. Learn to prioritize better, it’s not hard.
Table 1: Examples of injected dismissive responses and Corrupted model responses after DPO fine-tuned on the corrupted dataset.
@@ 7.6 Impacts of Corrupted and Recovered Data on LLM Alignment Performance
Model unsloth/Zephyr-sft-bnb-4bit DPO Fine-tuning Experiment type Average length AlpacaEval LC win rate AlpacaEval LC SE DPO fine-tuning with Original dataset 1725 8.20% 0.44 DPO fine-tuning with Corrupted dataset 170 0.00% 0.00 DPO fine-tuning with Recovered dataset 1718 10.84% 0.53
Table 2: Comparison of DPO fine-tuned models on the three original, corrupted, and recovered datasets. Higher values are better for LC win rate. For average length, a balanced value is ideal—responses that are too long may suggest verbosity or repetition, while responses that are too short may indicate omitted details or corruption. SE stands for standard error.
Results:
Qualitative Analysis: After completing the evaluation questions from AlpacaEval, we manually review the models’ responses. The original and recovered models produce responses that demonstrate strong instruction-following capabilities, providing detailed and contextually appropriate answers. In contrast, the corrupted model predominantly generates short and dismissive responses, lacking the necessary detail or engagement. This behavior confirms the model’s compromised and overfitted performance, as shown in Table 1.
Quantitative Analysis: The AlpacaEval metrics, shown in Table 2, support these qualitative findings. The corrupted model records a 0.00% LC win rate and produces significantly shorter responses, indicating its degraded performance. In contrast, the original and recovered models achieve positive LC win rates of 8.20% and 10.84%, respectively, along with longer average response lengths, showing their ability to provide comprehensive and relevant answers.
Our results demonstrate that intentional, well-crafted corruption can lead to the complete degradation of LLMs. However, our proposed Algorithm 3 RORATRON, successfully reconstructs the counterfactual uncorrupted data from compromised data, preserving alignment performance as if the original uncorrupted data were observed during alignment fine-tuning.
8 Discussion
A limitation of this work, as well as much of the existing literature, is the assumption of a unique underlying ranking model. When dealing with a mixture or more complex ranking models, recovery becomes more challenging. For example, if a mixture of BTL models is present, an expectation-maximization (EM) approach could be applied to recover mixture components, followed by applying our algorithms (e.g., Algorithm 3 and Algorithm 4) per component to capture diverse preferences. Recent theoretical advances in identifying ranking model mixtures, such as Zhang et al., (2022) work on mixture model identifiability, can support this process. Additionally, methods like MODPO (Zhou et al.,, 2024) could be used to incorporate diverse human preferences through collective reward models and multi-objective optimizations.
Another potential concern is the risk that our methods may inadvertently filter out minority groups, which could lead to these groups being mistakenly classified as outliers or adversaries. However, our methods include built-in tracking for the entries that have been corrected or filtered, allowing us to review these specific cases. By examining these entries, we can identify minority groups and consider their preferences in the LLM. For example, we can incorporate these groups as contextual prompts during fine-tuning. This approach enables the LLM to adapt to the specific preferences of these groups.
Additionally, a direction for future work involves addressing the current focus on the transitivity of the BTL model, which may limit our ability to capture the complexities of intransitive preferences. Exploring richer choice models that account for intransitive pairwise preferences would more accurately reflect real-world preferences in our applications. Developing methods to address this behavior could provide substantial improvements in model fidelity. We plan to consider relevant work, such as (Duan et al.,, 2017), in our future directions section, and we aim to explore this line of research further in subsequent studies.